# TRENDY: Gene Regulatory Network Inference Enhanced by Transformer

### v1

@@ 3 Discussion
In this paper, we present the TRENDY method, which uses the transformer model $\mathrm{TE}(k)$ to enhance the mechanism-based GRN inference method WENDY. TRENDY is tested against three other transformer-enhanced methods and their original forms, and ranks the first. This work explores the potential of deep learning methods in GRN inference when there are sufficient training data. Besides, TRENDY is developed on the biological dynamics of gene regulation, and it is more interpretable than other deep learning-based methods that treat GRN inference as a pure prediction task.
@@
The essential difficulty of GRN inference is the lack of data with experimentally verified GRNs. Many deep learning models need significantly more data with known GRNs, and we have to simulate new data from a generator. Since we do not fully understand the relation between the dynamics of gene expression and the ground truth GRN, it is difficult to evaluate the quality of generated data along with a random GRN. We also cannot guarantee that GRN inference methods that perform well on known experimental data sets can still be advantageous if there are many more experimental data with different GRNs.
@@
We train the models with data of $n=10$ genes, and they work well on data with $n=10/18/20$ genes. If we want the models to work on data with many more genes, we need to train the models with corresponding data. However, when the gene number $n$ is large, the number of different GRNs increases exponentially, and the number of training samples should also increase violently. The time cost for generating such data can be quite long. Even with the same amount of training samples, the training speed of the $\mathrm{TE}(k)$ model is proportional to $n^{4}$, since the input length of the transformer encoder layer is proportional to $n^{2}$, and it needs to calculate the attention scores between all pairs of input tokens. Therefore, scaling TRENDY for a much larger $n$ is extremely time-consuming.
@@
In this paper, the training data are generated by Eq. 1, which only describes the dynamics of mRNA count as a continuous-state process. This might be an oversimplification of the gene regulation in reality, where the gene state and the protein count also should be considered [42]. Besides, the experimental gene expression data often suffer from incomplete measurement, where many mRNAs are not recorded, and there are many missing values. Therefore, when we train on perfectly measured data and test on data with missing values, the performance is not guaranteed. We can manually add measurement errors and missing values to the training data to solve this problem.
In our testing, we find that using $\mathrm{TE}(k=1)$ model to directly enhance an inferred GRN, like SINCERITIES-rev or NonlinearODEs-rev, may not be sufficiently stable. Instead, using $\mathrm{TE}(k>1)$ model with both inferred GRN and corresponding covariance matrices, like TRENDY or GENIE3-rev, can achieve better results, since the covariance matrices may contain extra information besides that contained in the inferred GRN.
Some deep learning-based GRN inference methods directly use the raw gene expression data as the model input. We think that the covariance matrix contains most of the information regarding the relation between any two genes, and the covariance matrix is calculated from graphical lasso, thus being more regular than the original expression data. Therefore, using covariance matrices in deep learning GRN inference methods should be a promising approach.
@@ 4.1 Training data
To generate training data for our TRENDY method, we use Eq. 1 to numerically simulate gene expression data with the Euler-Maruyama method [43] for $n=10$ genes. Every time we generate a random ground truth GRN $A_{\mathrm{true}}$, where each element has probability $0.1/0.8/0.1$ to be $-1/0/1$. Then we use Eq. 1 with this random $A_{\mathrm{true}}$ to run the simulation from time $0.0$ to time $1.0$ with time step $0.01$ to obtain one trajectory, where the initial state at time 00 is random. This is repeated 100 times to obtain 100 trajectories, similar to the situation where we measure the single-cell expression levels of $n$ genes for 100 cells [44]. We only record expression levels at time points $0.0,0.1,0.2,\ldots,1.0$, and then calculate the covariance matrix for $n=10$ genes at each time point with graphical lasso.
We repeat the simulation to obtain $1.01\times 10^{5}$ samples, where $10^{5}$ samples are for training, and $10^{3}$ samples are for validation (used in hyperparameter tuning and early stopping). Since the TRENDY method only needs two time points, we fix the first time point to be $0.0$, and the second time point can be $0.1,0.2,\ldots,1.0$. Therefore, each simulated sample corresponds to 10 pairs of $K_{0}$ and $K_{t}$. This means that the training set has $10^{6}$ pairs of $K_{0}$ and $K_{t}$, where each $t$ in $0.1,0.2,\ldots,1.0$ corresponds to $10^{5}$ pairs.
For GENIE3-rev, we consider each time point $t$ in $0.1,0.2,\ldots,1.0$ and calculate the inferred GRN $A_{\mathrm{G}}$. Then there are $10^{6}$ pairs of $A_{\mathrm{G}}$ and $K_{t}$ for training the $\mathrm{TE}(k=2)$ model. For SINCERITIES-rev and NonlinearODEs-rev, we use the data at all time points $0.0,0.1,0.2,\ldots,1.0$ to infer the GRN $A_{\mathrm{S}}$ and $A_{\mathrm{N}}$. Therefore, each method has $10^{5}$ samples for training the $\mathrm{TE}(k=1)$ model.
@@ 4.2 Loss function
We use AUROC and AUPRC to measure the difference between $A_{\mathrm{true}}$ and the inferred $A_{\mathrm{pred}}$. Since $A_{\mathrm{true}}$ can take three values $-1/0/1$, the calculation of AUROC and AUPRC need some revisions [13]. Unfortunately, for fixed $A_{\mathrm{true}}$ and arbitrary $A_{\mathrm{pred}}$, AUROC and AUPRC can only take finitely many values, and are not continuous functions of $A_{\mathrm{pred}}$. We cannot train a neural network with a non-continuous loss function, since it is not differentiable, and the back propagation cannot work. Therefore, we cannot directly use AUROC or AUPRC as the loss function when training $\mathrm{TE}(k)$ models. There have been some surrogate loss functions for AUROC or AUPRC [45, 46], but we find that with enough training data, the mean square error (MSE) function can lead to satisfactory training results. We use MSE as the loss function for the $\mathrm{TE}(k=1)$ model in TRENDY. For all other $\mathrm{TE}(k)$ models used, the loss function is also MSE, but the diagonal elements are ignored, since they represent the autoregulation of one gene $i$ to itself, which cannot be inferred by the methods discussed in this paper [47].
@@ 4.3 Segment embedding
For $\mathrm{TE}(k)$ model with $k>1$ input matrices, after processing them with the linear embedding layer, we need to add segment embeddings to them. We give the segment ID $0/1/2/\ldots$ to the first/second/third/$\ldots$ group of processed inputs, and use an embedding layer to map each segment ID to a $d$-dimensional embedding vector. Repeat this embedding vector $n\times n$ times to obtain an $(n\times n\times d)$ array, and add it to the $(n\times n\times d)$ arrays after the linear embedding layer.
@@ 4.4 Positional encoding
For the $\mathrm{TE}(k)$ model, the 2-dimensional positional encoding layer [27] adds an $(n\times n\times d)$ array PE to each of the embedded input. For $x$ and $y$ in $1,2,\ldots,n$ and $j$ in $41,2,\ldots,d/$ 4, PE is defined as

### v2

@@ 3 Discussion
In this paper, we present the TRENDY method, which uses the transformer model $\mathrm{TE}(k)$ to enhance the mechanism-based GRN inference method WENDY. TRENDY is tested against three other transformer-enhanced methods and their original forms, and ranks the first. This work explores the potential of deep learning methods in GRN inference when there are sufficient training data. Besides, TRENDY is developed on the biological dynamics of gene regulation, and it is more interpretable than other deep learning-based methods that treat GRN inference as a pure prediction task. Nevertheless, the transformer encoder is directly used as an oracle machine in TRENDY. A future direction is to modify the neural network structure to better represent the biological structure [55].
@@
The essential difficulty of GRN inference is the lack of data with experimentally verified GRNs. Many deep learning models need significantly more data with known GRNs, and we have to generate new data. Since we do not fully understand the relation between the dynamics of gene expression and the ground truth GRN, it is difficult to evaluate the quality of generated data along with a random GRN. We also cannot guarantee that GRN inference methods that perform well on known experimental data sets can still be advantageous if there are many more experimental data with different GRNs.
@@
We train the models with data of $n=10$ genes, and they work well on data with $n=10/18/20$ genes. If we want the models to work on data with many more genes, we need to train the models with corresponding data. However, when the gene number $n$ is large, the number of different GRNs increases exponentially, and the number of training samples should also increase violently. The time cost for generating such data can be quite large. Even with the same amount of training samples, the training speed of the $\mathrm{TE}(k)$ model is proportional to $n^{4}$, since the input length of the transformer encoder layer is proportional to $n^{2}$, and it needs to calculate the attention scores between all pairs of input tokens. Therefore, scaling TRENDY for a much larger $n$ is extremely time-consuming.
@@
In this paper, the training data are generated by Eq. 1 that simulates the dynamics of mRNA count as a continuous-state process. This might be an oversimplification of gene regulation in reality, where the gene state and the protein count also should be considered [48]. Besides, the experimental gene expression data often suffer from incomplete measurement, where many mRNAs are not recorded, and there are many missing values. Therefore, when we train on perfectly measured data and test on data with missing values, the performance is not guaranteed. We can manually add measurement errors and missing values to the training data to solve this problem.
In our testing, we find that adding covariance matrices in the transformer (TRENDY and tGENIE3) is beneficial, since the covariance matrices may contain extra information besides that contained in the inferred GRN.
TRENDY uses covariance matrices instead of raw gene expression data as input. One advantage is that calculating covariance matrices by graphical lasso regularizes the data. One disadvantage is that covariance only concerns second-order moment, while information contained in higher-order moments is abandoned.
@@ 4.1 Training data
To generate training data for our TRENDY method, we use Eq. 1 to numerically simulate gene expression data with the Euler-Maruyama method [56] for $n=10$ genes. Every time we generate a random ground truth GRN $A_{\mathrm{true}}$, where each element has probability $0.1/0.8/0.1$ to be $-1/0/1$. Then we use Eq. 1 with this random $A_{\mathrm{true}}$ to run the simulation from time $0.0$ to time $1.0$ with time step $0.01$ to obtain one trajectory, where the initial state at time 00 is random. This is repeated 100 times to obtain 100 trajectories, similar to the situation where we measure the single-cell expression levels of $n$ genes for 100 cells [57]. We only record expression levels at time points $0.0,0.1,0.2,\ldots,1.0$. We repeat the simulation to obtain $1.01\times 10^{5}$ samples, where $10^{5}$ samples are for training, and $10^{3}$ samples are for validation (used in hyperparameter tuning and early stopping).
@@ 4.2 Loss function
We use AUROC and AUPRC to measure the difference between $A_{\mathrm{true}}$ and the inferred $A_{\mathrm{pred}}$. See Section S2 for details. Unfortunately, for fixed $A_{\mathrm{true}}$ and arbitrary $A_{\mathrm{pred}}$, AUROC and AUPRC can only take finitely many values, and are not continuous functions of $A_{\mathrm{pred}}$. We cannot train a neural network with a non-differentiable loss function like AUROC or AUPRC. There have been some surrogate loss functions for AUROC or AUPRC [58, 59]. We directly use the mean square error loss function, which also performs well. The loss for comparing two GRNs does not count the diagonal elements, since they represent the autoregulation that cannot be inferred by the methods in this paper [60].
@@ 4.3 Segment embedding
For $\mathrm{TE}(k)$ model with $k>1$ input matrices, after processing them with the linear embedding layer, we need to add segment embeddings to them, so that the model can differentiate between different input matrices [61]. We give the segment ID $0/1/2/\ldots$ to the first/second/third/$\ldots$ group of processed inputs, and use an embedding layer to map each segment ID to a $d$-dimensional embedding vector. Repeat this embedding vector $n\times n$ times to obtain an $(n\times n\times d)$ array, and add it to the $(n\times n\times d)$ arrays after the linear embedding layer. In our numerical simulations, we find that this segment embedding layer slightly improves the performance.
@@ 4.4 Positional encoding
For the $\mathrm{TE}(k)$ model, the 2-dimensional positional encoding layer [29] adds an $(n\times n\times d)$ array PE to each of the embedded input, so that the model knows that the input is two-dimensional. For $x$ and $y$ in $1,2,\ldots,n$ and $j$ in $41,2,\ldots,d/$ 4, PE is defined as
@@ 4.5 Training setting and cost
Appendix S1 Details of WENDY method
We start with a general SDE for gene expression levels $X_{j}(t)$:
$\frac{\mathrm{d}X_{i}(t)}{\mathrm{d}t}=f(X_{1},\ldots,X_{n})+c_{1,i}-c_{2,i}X_{i}+\mathrm{noise}.$
Here $f$ describes the interactions of different genes; $c_{1,i}$ is the synthesis rate; $c_{2,i}$ is the degradation rate.
After some simplifications and linearizations, we have
$\frac{\mathrm{d}X(t)}{\mathrm{d}t}=X(t)A+c+X(t)\odot\mathrm{d}\sigma W(t).$
Here $X(t)=[X_{1}(t),\ldots,X_{n}(t)]$, $A$ is the GRN, $W(t)$ is an $n$-dimensional standard Brownian motion, and $\odot$ is the entrywise (Hadamard) product.
Its approximated solution is
$X(t)=X(0)(I+tA)+tc+X(0)\odot\epsilon(t),$
where $\epsilon(t)=[\epsilon_{1}(t),\ldots,\epsilon_{n}(t)]$ is an $n$-dimensional normal random noise with 00 mean and diagonal covariance matrix.
Then we can calculate the covariance matrix
$K(t)=\mathbb{E}\{[X(t)^{\mathrm{T}}-\mathbb{E}X(t)^{\mathrm{T}}][X(t)-\mathbb{E}X(t)]\},$
and the final equation is
$K(t)=(I+tA^{\mathrm{T}})K(0)(I+tA)+D+E,$
where $D$ is an unknown diagonal matrix that depends on the variance of $\epsilon$, and $E$ is the error introduced by linearizations.
WENDY method solves $A$ from the last equation by the BFGS algorithm.
Appendix S2 Calculation of AUROC and AUPRC
In our setting, each edge in the true GRN can only take three values (labels): -1 (negative regulation), 0 (no regulation), and 1 (positive regulation). In the inferred GRN, each edge corresponds to a real number that can be mapped to those three labels. We need to evaluate the inferred GRN as a classification problem.
AUROC (Area Under the Receiver Operating Characteristic Curve) and AUPRC (Area Under the Precision-Recall Curve) are metrics for evaluating the performance of classification models (originally for binary classification). They are both between 0 and 1. If the case has perfect match, where the sorted predicted results and the sorted true labels have the same order, then AUROC and AUPRC are 1. If the order is fully reversed, AUROC and AUPRC are 0. Notice that only the order matters, and the predicted results do not need to match the true labels numerically. Although they have similar properties, AUROC is more valuable when different true labels are relatively balanced, and AUPRC is more informative with imbalanced labels. Therefore, we use both to evaluate the inferred GRN.
We use a toy example to illustrate the calculation of AUROC and AUPRC, given the ground truth GRN and the inferred GRN.
Assume that the ground truth GRN and the inferred GRN are
$A_{\mathrm{true}}=\begin{bmatrix}0&0&0&1&1\\ 1&1&-1&0&0\\ 0&-1&1&0&1\\ 1&-1&-1&1&0\\ 0&-1&-1&-1&-1\end{bmatrix},$
and
$A_{\mathrm{pred}}=\begin{bmatrix}0.45&0.15&0.30&-0.94&0.24\\ 0.02&0.09&-0.28&-0.84&-0.61\\ 0.90&-0.18&-0.07&-0.89&-0.92\\ 0.43&0.08&0.02&-0.51&0.49\\ -0.66&-0.62&-0.08&-0.43&-0.51\end{bmatrix}.$
Since all GRN inference methods in the main text cannot handle autoregulation, we omit edges $(i,i)$. For each edge $(i,j)$ with $i\neq j$, we know the true label $A_{\mathrm{true}}[i,j]$ and the predicted value $A_{\mathrm{pred}}[i,j]$. Since the true label has three possibilities, but AUROC and AUPRC are originally defined for binary classification, we need to consider every two different labels.
Consider all edges $(i,j)$ with $i\neq j$ and true label 00 or $1$. List $(A_{\mathrm{pred}}[i,j],A_{\mathrm{true}}[i,j])$ and sort by $A_{\mathrm{pred}}[i,j]$. Then we have (-0.94, 1), (-0.92, 1), (-0.89, 0), (-0.84, 0), (-0.66, 0), (-0.61, 0), ( 0.02, 1), ( 0.15, 0), ( 0.24, 1), ( 0.30, 0), ( 0.43, 1), ( 0.49, 0), ( 0.90, 0).
For a given threshold $T$, we can count
true positives (TP): $A_{\mathrm{true}}[i,j]=1$, $A_{\mathrm{pred}}[i,j]>T$;
true negatives (TN): $A_{\mathrm{true}}[i,j]=0$, $A_{\mathrm{pred}}[i,j]\leq T$;
false positives (FP): $A_{\mathrm{true}}[i,j]=0$, $A_{\mathrm{pred}}[i,j]>T$;
false negatives (FN): $A_{\mathrm{true}}[i,j]=1$, $A_{\mathrm{pred}}[i,j]\leq T$.
Then we can calculate
true positive rate (TPR): TP/(TP+FN);
false positive rate (FPR): FP/(FP+TN).
For instance, if the threshold to be 00, then
TP = 3,  TN = 4,  FP = 4,  FN = 2,  TPR = 0.6,  FPR = 0.5.
We let the threshold $T$ change from -1 to 1, and draw the corresponding points (FPR, TPR) in the square $[0,1]\times[0,1]$: (0.0, 0.0), (0.125, 0.0), (0.25, 0.0), (0.25, 0.2), (0.375, 0.2), (0.375, 0.4), (0.5, 0.4), (0.5, 0.6), (1.0, 0.6), (1.0, 1.0). Connect these points to obtain the receiver operating characteristic (ROC) curve, and AUROC is the area under this curve. See Fig. S1 for the illustration, where AUROC is 0.375.
Figure S1: Example of ROC curve, where AUROC is the shaded area.
For true labels -1 and 1, repeat this procedure, and the AUROC[-1,1] is 0.571. For true labels -1 and 0, repeat this procedure, and the AUROC[-1,0] is 0.518.
AUROC for two labels has an equivalent definition: in all pairs of edges $(i,j)$ with $i\neq j$ and $(p,q)$ with $p\neq q$, so that $A_{\mathrm{true}}[i,j]<A_{\mathrm{true}}[p,q]$, AUROC is the proportion that $A_{\mathrm{pred}}[i,j]<A_{\mathrm{pred}}[p,q]$.
Inspired by this definition, the AUROC for three labels should be the weighted average of three AUROC for two labels, where the weight is the number of pairs with different labels. In this example, there are 5 edges with true label 1, 8 edges with true label 0, and 7 edges with true label -1. Therefore, the final AUROC is
$\frac{5\times 8\times\mathrm{AUROC}[0,1]+5\times 7\times\mathrm{AUROC}[-1,1]+8\times 7\times\mathrm{AUROC}[-1,0]}{5\times 8+5\times 7+8\times 7}=0.489.$
To calculate AUPRC, just replace TPR by Precision=TP/(TP+FP), and replace FPR by Recall=TP/(TP+FN).
Appendix S3 Training and testing details
During training, for variants of WENDY, since they only need two time points, we fix the first time point to be $0.0$, and the second time point can be $0.1,0.2,\ldots,1.0$. Therefore, each simulated sample corresponds to 10 pairs of $K_{0}$ and $K_{t}$. This means that the training set has $10^{6}$ pairs of $K_{0}$ and $K_{t}$, where each $t$ in $0.1,0.2,\ldots,1.0$ corresponds to $10^{5}$ pairs. For variants of GENIE3, we consider each time point $t$ in $0.1,0.2,\ldots,1.0$ and calculate the inferred GRN $A_{\mathrm{G}}$. Then there are $10^{6}$ pairs of $A_{\mathrm{G}}$ and $K_{t}$ for training the $\mathrm{TE}(k=2)$ model. For variants of SINCERITIES and NonlinearODEs, we use the data at all time points $0.0,0.1,0.2,\ldots,1.0$ to infer the GRN $A_{\mathrm{S}}$ and $A_{\mathrm{N}}$. Therefore, each method has $10^{5}$ samples for training the $\mathrm{TE}(k=1)$ model.
For SINC data set, there are 1000 samples for each value of $\sigma=0.01/0.1/1$, and each sample with a random $A_{\mathrm{true}}$ has 100 cells measured at time points $0.0,0.1,\ldots,1.0$. For WENDY and its variants, we set the first time point to $0.0$ and vary the second time point from $0.1$ to $1.0$. For GENIE3 and its variants, we set the time point to any of $0.1,\ldots,1.0$. For SINCERITIES, NonlinearODEs, and their variants, we use all time points.
For DREAM4 data set, we use the 10-gene time series data set. There are five $A_{\mathrm{true}}$, each corresponding to the levels of 10 genes for five cells at 21 time points. For WENDY and its variants, we consider any pair of time points and take average. For GENIE3 and its variants, we consider any one time point and take average. For SINCERITIES, NonlinearODEs, and their variants, we use any consecutive 11 time points and take average.
For the THP-1 data set, there are eight time points, each with 120 cells measured. We consider the expression levels of 20 genes. See Fig. S2 for the ground truth GRN. For WENDY and its variants, we consider any pair of time points and take average. For GENIE3 and its variants, we consider any one time point and take average. For SINCERITIES, NonlinearODEs, and their variants, we use all eight time points.
For the hESC data set, there are six time points, each with 66–172 cells measured. The same as in a previous paper [15], we consider the expression levels of 18 genes. See Fig. S3 for the ground truth GRN. For WENDY and its variants, we consider any pair of time points and take average. For GENIE3 and its variants, we consider any one time point and take average. For SINCERITIES, NonlinearODEs, and their variants, we use all six time points.
Appendix S4 Heat maps of true GRN for experimental data
See Fig. S2 for the heat map of the true GRN of THP-1 data set. Possible values are 1, -1, and 0, meaning positive regulation, negative regulation, and no regulation.
Figure S2: Heat map of the true GRN of THP-1 data set.
See Fig. S3 for the heat map of the true GRN of hESC data set. Possible values are 1 and 0, meaning regulation and no regulation. Notice that almost all values are 0.
Figure S3: Heat map of the true GRN of hESC data set.
Appendix S5 Detailed performance measurements
$\sigma=0.01$ $\sigma=0.1$ $\sigma=1$ Average
Method AUROC AUPRC AUROC AUPRC AUROC AUPRC AUROC AUPRC
WENDY 0.6290 0.5897 0.6654 0.6109 0.6008 0.5782 0.6317 0.5929
TRENDY 0.7533 0.6908 0.8547 0.7519 0.6678 0.6087 0.7586 0.6838
nWENDY 0.5227 0.5431 0.5131 0.5447 0.5045 0.5314 0.5134 0.5397
bWENDY 0.5227 0.5413 0.5129 0.5431 0.5052 0.5312 0.5136 0.5385
GENIE3 0.4487 0.5378 0.4062 0.5061 0.3965 0.4881 0.4171 0.5107
tGENIE3 0.7844 0.7189 0.8703 0.7672 0.6757 0.6040 0.7768 0.6967
nGENIE3 0.4403 0.5241 0.3985 0.4976 0.3689 0.4748 0.4026 0.4988
bGENIE3 0.4418 0.5223 0.3983 0.4959 0.3693 0.4746 0.4031 0.4976
SINCERITIES 0.6493 0.5726 0.6783 0.5829 0.7154 0.5967 0.6810 0.5841
tSINCERITIES 0.7192 0.6185 0.7964 0.6637 0.7661 0.6294 0.7606 0.6372
nSINCERITIES 0.5401 0.5369 0.5394 0.5387 0.5332 0.5363 0.5376 0.5373
bSINCERITIES 0.5422 0.5375 0.5450 0.5409 0.5374 0.5381 0.5415 0.5388
NonlinearODEs 0.5026 0.5243 0.5076 0.5313 0.5068 0.5301 0.5057 0.5286
tNonlinearODEs 0.5487 0.5460 0.5976 0.5658 0.5162 0.5333 0.5542 0.5484
nNonlinearODEs 0.4971 0.5205 0.5053 0.5251 0.5073 0.5236 0.5032 0.5231
bNonlinearODEs 0.4972 0.5202 0.5058 0.5260 0.5081 0.5248 0.5037 0.5237
Table S1: AUROC and AUPRC scores of all 16 methods on SINC data set. We list the scores separately for data with $\sigma=0.01$, $\sigma=0.1$, $\sigma=1$, and the average scores.
DREAM4 THP-1 hESC
Method AUROC AUPRC AUROC AUPRC AUROC AUPRC
WENDY 0.4899 0.2080 0.5261 0.3972 0.4997 0.0392
TRENDY 0.5341 0.2177 0.5557 0.3669 0.5311 0.0376
nWENDY 0.5417 0.2254 0.6112 0.4203 0.4971 0.0372
bWENDY 0.5421 0.2231 0.6106 0.4205 0.5070 0.0402
GENIE3 0.5636 0.2286 0.4484 0.3546 0.5913 0.0468
tGENIE3 0.4589 0.1799 0.5506 0.3781 0.6008 0.0435
nGENIE3 0.5632 0.2261 0.4861 0.3642 0.5744 0.0462
bGENIE3 0.5741 0.2284 0.4792 0.3623 0.5767 0.0488
SINCERITIES 0.4908 0.1919 0.6261 0.3852 0.4198 0.0261
tSINCERITIES 0.4995 0.2034 0.5251 0.3412 0.4871 0.0294
nSINCERITIES 0.4999 0.1856 0.5956 0.3900 0.1955 0.0199
bSINCERITIES 0.5040 0.1846 0.6067 0.3798 0.1842 0.0196
NonlinearODEs 0.4806 0.1705 0.5338 0.3486 0.5971 0.0534
tNonlinearODEs 0.5712 0.2452 0.4808 0.3302 0.6233 0.0641
nNonlinearODEs 0.4791 0.1772 0.5521 0.3482 0.6008 0.0466
bNonlinearODEs 0.4856 0.1666 0.5544 0.3498 0.6040 0.0633
Table S2: AUROC and AUPRC scores of all 16 methods on DREAM4, THP-1, and hESC data sets.